Calculate the molar Gibbs free energy (in kJ) of mixing 5moles of He(g) with 10moles of H2(g). The separate gases and the gas mixture are all at a pressure of 1 bar and at a temperature of 36oC. Assume that both the separate gases and the mixture behave ideally.
1
Expert's answer
2012-04-03T11:28:20-0400
dG = -TdSdS = k(n1*ln(n/n1) + n2*ln(n/n2))
n=n1 + n2
k-Boltzman constant = 1.38×10^-23 J/K

dS = 1.38×10^-23 (5*ln15/5 + 10*ln15/10) = 1.318 x10^-22

dG = -(273 + 36)*1.318x10^-22 = 4.07x10^-20 J = 4.07x10^-17 k